OpenStudy (anonymous):

when a polynomial f(x)divided by (x-1)it gives remainder 4 &when divided by (x-2) remainder=5 what will be remainder if f(x)is dividedby (x-1)(x-2)

OpenStudy (anonymous):

love remainder/factor theorem

OpenStudy (anonymous):

P(x)=Q(x)D(x)+R

OpenStudy (amistre64):

\[\frac{x-1+4}{x-1}\] \[\frac{x-2+5}{x-2}\] since f(x) = x+3 is "a" solution .... i wonder if we can be this specific

OpenStudy (anonymous):

i think we look more at the remainder theorem here.

OpenStudy (anonymous):

For this speicfic question P(x)=Q(x)(x-1)(x-2)+(Ax+B)

OpenStudy (anonymous):

The remainder is a linear factor as it is 1 degree lower than a quadratic

OpenStudy (anonymous):

Is this guy online?

OpenStudy (amistre64):

prolly not .... but the question was nice

OpenStudy (anonymous):

Anyway we know P(1)=4 from first lot of info and P(2)=5

OpenStudy (anonymous):

now its just simultaneous equations

OpenStudy (anonymous):

So we get A+B=4 2A+B=5

OpenStudy (anonymous):

A=1 B=3

OpenStudy (anonymous):

tHEREFORE remainder is x+3

OpenStudy (amistre64):

so i was right :)

OpenStudy (anonymous):

which is what u got i guess haha

OpenStudy (anonymous):

i gotta hand it to you